OpenStudy (anonymous):

MEDAL You are going to create a circle graph to represent some data. How many degrees should a section representing 65 out of a hundred be? 234° 65° 100° 13°

OpenStudy (anonymous):

@lopezking1 @chmvijay @Nurali @robtobey @satellite73

OpenStudy (lopezking1):

How many degrees are in a circle

OpenStudy (chmvijay):

65/100= x /degrees in circle How many degrees are in a full circle ?

OpenStudy (nurali):

65/100= x /degrees in circle 65/100= x /360 65*360/100=x x=234

OpenStudy (chmvijay):

yaa that right =360 °C

OpenStudy (anonymous):

sorry...didn't saw itt